# Do Not Design, Learn: A Trainable Scoring Function for Uncertainty Estimation in Generative LLMs

### v1

@@ 5.5 LARS without Labeled Data
6 Ablation Studies
6.1 Probability Association Strategies
In Section 4, we explain a sequential approach to associate tokens of the response with their probabilities, where special probability tokens are placed after each response token in the input to LARS. As an alternative, we explore an additive approach. In this method, the embedding vectors of the probability tokens are added to the embedding vectors of their corresponding response tokens. This strategy effectively reduces the input sequence length for the LARS model. Results (see Figure 3) demonstrate that the sequential approach is, on average, 0.15 points better when used with Confidence, although the gap narrows for Entropy and SE. Comparing the additive approach with other baselines from Table 1, we observe that it still significantly outperforms the baselines. Overall, these two probability association approaches highlight a possible trade-off between shortened input length (to the LARS model) and improved UE performance.
6.2 Size of the Calibration Dataset
To evaluate the scalability of LARS, we calibrate it using different amounts of labeled data. The results, depicted in Figure 4, show that even with as few as 1,000 labeled question-ground truth pairs, LARS outperforms the best-performing baseline. More notably, LARS demonstrates good scalability with calibration data size. Exploring the scaling of LARS with even more data remains as a future direction.
6.3 Importance of LARS Input Components
In this section, we assess the impact of individual input components of LARS on UE.
Number of Probability Tokens.
@@ 6.3 Importance of LARS Input Components
Effect of Probability Information.
To assess the importance of probability information for LARS, we train a version of the model using only textual inputs: the question and the generated answer. The results (Table 6) indicate that excluding probability information leads to a decrease in the performance of LARS by up to 0.101 points. This significant drop underscores the critical role that probability information plays in the efficacy of LARS.
Effect of Textual Information.
UE Method Scoring Function AUROC
Confidence Only text 0.750
Only probs 0.721
LARS 0.851
Entropy Only text 0.754
Only probs 0.733
LARS 0.842
SE Only text 0.817
Only probs 0.799
LARS 0.849
Table 6: Comparison of AUROC performance for the Llama2-7b model on the TriviaQA Dataset across different input modalities: text-only, probabilities-only, and combined text and probabilities.
7 Conclusion
In this study, we demonstrated the shortcomings of existing scoring functions and introduced LARS, an off-the-shelf scoring function directly learned from data. We demonstrated that LARS significantly outperforms existing baselines across three different QA datasets with low computational cost. Additionally, we showed that LARS can be effectively trained even without labeled data, by using a teacher labeling model, and still surpasses the performance of the teacher model. Furthermore, our results indicate that LARS’ performance scales well with increased data.
8 Limitations
One limitation of LARS is its reliance on labeled data, which is not a requirement for other scoring functions. While LARS shows promise in environments without labeled data, this aspect requires further investigation to enhance its performance. Further, LARS depends on a pretrained RoBERTa model, which has a limited sequence length capability. This may necessitate the pre-training of Bert-like models that can handle longer sequences. Lastly, training LARS with a transformer model reduces the interpretability of the features. Traditional scoring functions modify the weighting of probabilities and compute a dot product between log probabilities and weights, offering a level of interpretability that LARS, with its more complex function (despite its superior performance), lacks.
9 Ethics Statement
Appendix A Details of Turkish Experiment
We translate the same 13k question-ground truth pairs from the train split of TriviaQA to Turkish using Googletrans library [^2] https://py-googletrans.readthedocs.io. Then, we apply the same procedure as for English: Make the LLM generate 6 answers to the question, ensuring the most likely generation is included. To train LARS, we utilize unique question-response pairs. The labels for training are again obtained by using GPT-3.5-turbo for each QA pair.
To test the performance of varying scoring functions in Turkish, we also translate the question-ground truth pairs of the same test samples of TriviaQA. The same multiple-generation procedure is performed for this set as well and the label is obtained for the most likely generation. After having the translated test set, the Entropy UE metric is calculated by using various scoring functions.
Lastly, the prompts for the LLM are also translated into Turkish to make sure it provides answers in Turkish. Prompts are provided below.
For Llama3-8b to generate answers:  [^3] English translation: You are a helpful, respectful and honest assistant. Give short and precise answers to given questions.
System: Sen yardmc, saygl ve drst
bir asistansn. Sorularm Trke olacak
ekilde net, ksa ve z cevapla.
User: {question}
For GPT-3.5-turbo to obtain labels:
You will behave as a question answer
evaluator. I will give you a question,
the ground truth of the question, and
a generated answer by a language model
in Turkish. You will output "correct"
if the generated answer is correct
regarding question and ground truth.
Otherwise, output "false".
Question: {question},
Ground Truth: {gt_answer},
Generated Answer: {generation}
Appendix B Details of LARS training
We use the pre-trained RoBERTa-base model with a single logit fully-connected layer added at the end. Binary cross entropy loss is used, while the optimizer is AdamW with a learning rate of $65e-$ 6. The model is trained for 5 epochs. We did a search for batch size in the set of $\{4,8,16,32\}$ and found the optimal batch size as 8 and used it in all of the experiments. The search set for learning rate was $\{1e-6,5e-6,1e-5,5e-4,1e-4,5e-4\}$. Lastly, we explored training the model for more epochs (up to 10); however, after epoch 5, we observed overfitting.
Appendix C Additional Experiments
C.1 OOD Model Experiments - LARS
@@
C.2 OOD Data Experiments - LARS
UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
Confidence Best Score of Baselines 0.7137 0.7166 0.6923 0.6453
LARS (NaturalQA only) 0.7685 0.7940 0.7765 0.7846
LARS (TriviaQA only) 0.7455 0.7689 0.7365 0.7456
LARS (TriviaQA+NaturalQA) 0.7731 0.7997 0.7774 0.7818
Entropy Best Score of Baselines 0.7071 0.7144 0.7014 0.6459
LARS (NaturalQA only) 0.7655 0.7936 0.7781 0.7786
LARS (TriviaQA only) 0.7434 0.7736 0.7392 0.7468
LARS (TriviaQA+NaturalQA) 0.7629 0.7918 0.7761 0.7814
SE Best Score of Baselines 0.7301 0.7591 0.7352 0.6701
LARS (NaturalQA only) 0.7665 0.7873 0.7770 0.7758
LARS (TriviaQA only) 0.7511 0.7750 0.7497 0.7572
LARS (TriviaQA+NaturalQA) 0.7635 0.7849 0.7766 0.7760
Table 8: OOD data experiments on NaturalQA dataset
UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
Confidence Best Score of Baselines 0.7510 0.7994 0.7468 0.6043
LARS (TriviaQA only) 0.8505 0.8721 0.8443 0.8350
LARS (NaturalQA only) 0.7780 0.8243 0.7893 0.7720
LARS (TriviaQA+NaturalQA) 0.8414 0.8620 0.8305 0.8152
Entropy Best Score of Baselines 0.7356 0.8012 0.7634 0.6053
LARS (TriviaQA only) 0.8381 0.8514 0.8213 0.8415
LARS (NaturalQA only) 0.7852 0.8348 0.8090 0.7775
LARS (TriviaQA+NaturalQA) 0.8354 0.8602 0.8373 0.8145
SE Best Score of Baselines 0.7973 0.8451 0.8132 0.7318
LARS (TriviaQA only) 0.8488 0.8662 0.8541 0.8281
LARS (NaturalQA only) 0.8181 0.8515 0.8349 0.7911
LARS (TriviaQA+NaturalQA) 0.8457 0.8621 0.8493 0.8184
Table 9: OOD data Experiments on TriviaQA dataset
@@ C.2 OOD Data Experiments - LARS
Table 10: Data samples from the datasets we use to evaluate UE methods: TriviaQA, NaturalQA, and WebQA.
@@ Appendix D Experimental Details
Datasets. To train the LARS model, for each TriviaQA and NaturalQA training split, we randomly select $\sim$13k samples resulting in $\sim$60k sampled unique QA pairs. To evaluate the UE methods we use 3 datasets: $\sim$9k samples from the TriviaQA validation split, the validation set of NaturalQA consisting of $\sim$3500 samples, and $\sim$6k samples coming from the train and validation sets of WebQA combined.
@@
Computational Cost. We use 40 GB Nvidia A-100 GPUs for all the experiments. The total GPU-hours for training a LARS model with a calibration dataset generated from $\sim$13k questions is approximately 4. Labeling of the calibration data for one dataset and one model takes approximately 30 GPU-hours. Getting all the results in Table 1 compromises $\sim$230 GPU-hours excluding LARS training. All presented results are obtained with a single run.
@@
User: {question}
@@
The prompt for the teacher models explained in Section 5.5 is as follows:

### v2

@@ C.6 Choice of Encoder-only Transformer
UE Method Scoring English Turkish German Spanish
Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR
Lex. Sim. - 0.683 0.483 0.652 0.361 0.660 0.410 0.640 0.369
# Sem. Gr. - 0.819 0.671 0.683 0.343 0.774 0.571 0.787 0.605
Eccentricity - 0.853 0756. 0.732 0.482 0.805 0.668 0.813 0.689
Degree Matrix - 0.851 0.746 0.748 0.583 0.802 0.675 0.814 0.692
Confidence LNS 0.748 0.600 0.714 0.500 0.727 0.544 0.704 0.504
MARS 0.799 0.676 0.720 0.503 0.747 0.582 0.728 0.550
TokenSAR 0.792 0.674 0.747 0.568 0.779 0.645 0.761 0.609
LARS 0.872 0.817 0.831 0.703 0.843 0.754 0.852 0.764
Entropy LNS 0.747 0.594 0.692 0.450 0.710 0.523 0.701 0.491
MARS 0.801 0.672 0.695 0.457 0.728 0.555 0.723 0.533
TokenSAR 0.793 0.676 0.720 0.515 0.758 0.614 0.750 0.590
LARS 0.864 0.804 0.814 0.680 0.827 0.737 0.835 0.742
SE LNS 0.835 0.733 0.734 0.554 0.791 0.663 0.797 0.667
MARS 0.835 0.742 0.734 0.551 0.789 0.663 0.795 0.666
TokenSAR 0.839 0.747 0.739 0.568 0.796 0.676 0.800 0.677
LARS 0.866 0.811 0.799 0.668 0.824 0.735 0.831 0.742
SentSAR LNS 0.825 0.723 0.728 0.530 0.765 0.629 0.765 0.617
MARS 0.838 0.746 0.731 0.531 0.775 0.641 0.775 0.631
TokenSAR 0.840 0.750 0.752 0.577 0.793 0.673 0.790 0.660
LARS 0.879 0.823 0.828 0.705 0.841 0.757 0.848 0.763
Table 13: AUROC and PRR performance of different UE methods on Llama3-8B for the TriviaQA dataset in different languages.
C.7 Effect of Number of Probability Tokens
@@
C.8 LARS on Different Languages
Extended results on different languages from Section 5.4 are presented in Table 13. In all languages, LARS consistently outperforms baseline UE methods and scoring functions.
C.9 Impact of Question as LARS Input
In this section, we evaluate the impact of including the question as input to the LARS model. The results, shown in Table 14, indicate a consistent performance drop when the question is omitted. This outcome is expected, as the question context plays a crucial role in determining whether a generated response is nonsensical or off-topic. However, we argue that the performance drop is not substantial. Thus, if computational efficiency is a priority, LARS can still be effectively used without the question context.
UE Method Scoring Function AUROC PRR
Confidence No question 0.832 0.720
LARS 0.851 0.760
Entropy No question 0.828 0.716
LARS 0.842 0.748
SE No question 0.836 0.736
LARS 0.849 0.745
SentSAR No question 0.842 0.740
LARS 0.850 0.763
Table 14: LARS with and without question in the input.
C.10 OOD Model Experiments - LARS
@@
Dataset UE Method Calibration Model Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
TriviaQA Confidence Best Baseline Score 0.7510 0.7994 0.7468 0.6883 0.7278
Llama2-7b 0.8577 0.8355 0.8309 0.7993 0.8432
Llama3-8b 0.8519 0.8737 0.8499 0.7830 0.8146
Mistral-7b 0.8352 0.8327 0.8518 0.7872 0.8315
Gemma-7b 0.8169 0.8172 0.8097 0.8311 0.8054
Llama2-13b 0.8376 0.8526 0.8327 0.7829 0.8510
Entropy Best Baseline Score 0.7356 0.8012 0.7634 0.6941 0.7332
Llama2-7b 0.8469 0.8298 0.8271 0.8060 0.8473
Llama3-8b 0.8520 0.8730 0.8501 0.7955 0.8313
Mistral-7b 0.8410 0.8407 0.8542 0.7974 0.8380
Gemma-7b 0.8145 0.8181 0.8200 0.8279 0.8115
Llama2-13b 0.8335 0.8525 0.8435 0.7916 0.8553
SE Best Baseline Score 0.7973 0.8388 0.8132 0.7528 0.8062
Llama2-7b 0.8497 0.8358 0.8402 0.8100 0.8603
Llama3-8b 0.8625 0.8719 0.8623 0.8008 0.8518
Mistral-7b 0.8496 0.8490 0.8591 0.8056 0.8579
Gemma-7b 0.8334 0.8425 0.8372 0.8289 0.8386
Llama2-13b 0.8447 0.8660 0.8554 0.8049 0.8671
SentSAR Best Baseline Score 0.7940 0.8402 0.8050 0.7411 0.7910
Llama2-7b 0.8572 0.8409 0.8413 0.7900 0.8584
Llama3-8b 0.8663 0.8838 0.8634 0.8017 0.8453
Mistral-7b 0.8514 0.8474 0.8596 0.8055 0.8513
Gemma-7b 0.8309 0.8363 0.8327 0.8338 0.8272
Llama2-13b 0.8456 0.8687 0.8506 0.7980 0.8616
NaturalQA Confidence Best Baseline Score 0.7032 0.7136 0.6823 0.6433 0.6774
Llama2-7b 0.7886 0.7546 0.7512 0.7288 0.7613
Llama3-8b 0.7732 0.8113 0.7679 0.7265 0.7517
Mistral-7b 0.7538 0.7543 0.7868 0.7195 0.7507
Gemma-7b 0.7522 0.7397 0.7493 0.8033 0.7295
Llama2-13b 0.7727 0.7695 0.7571 0.7308 0.7812
Entropy Best Baseline Score 0.6831 0.7144 0.6944 0.6439 0.6859
Llama2-7b 0.7756 0.7582 0.7550 0.7367 0.7641
Llama3-8b 0.7734 0.8103 0.7767 0.7374 0.7544
Mistral-7b 0.7569 0.7642 0.7877 0.7305 0.7607
Gemma-7b 0.7481 0.7463 0.7506 0.7939 0.7351
Llama2-13b 0.7671 0.7752 0.7569 0.7363 0.7783
SE Best Baseline Score 0.7210 0.7591 0.7272 0.7049 0.7361
Llama2-7b 0.7695 0.7590 0.7574 0.7521 0.7716
Llama3-8b 0.7767 0.8038 0.7820 0.7513 0.7661
Mistral-7b 0.7627 0.7681 0.7826 0.7484 0.7680
Gemma-7b 0.7517 0.7602 0.7561 0.7916 0.7511
Llama2-13b 0.8049 0.7837 0.7672 0.7548 0.7843
SentSAR Best Baseline Score 0.7177 0.7563 0.7268 0.6891 0.7319
Llama2-7b 0.7835 0.7639 0.7595 0.7423 0.7738
Llama3-8b 0.7816 0.8154 0.7838 0.7417 0.7655
Mistral-7b 0.7669 0.7705 0.7940 0.7360 0.7682
Gemma-7b 0.7572 0.7567 0.7616 0.7978 0.7486
Llama2-13b 0.7980 0.7838 0.7654 0.7415 0.7853
Table 15: OOD model experiments on TriviaQA and NaturalQA datasets with AUROC scores.
@@
GSM8K Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? 72
Julie is reading a 120-page book. Yesterday, she was able to read 12 pages and today, she read twice as many pages as yesterday. If she wants to read half of the remaining pages tomorrow, how many pages should she read? 42
Mr. Sam shared a certain amount of money between his two sons, Ken and Tony. If Ken got \$1750, and Tony got twice as much as Ken, how much was the money shared? 5250
Table 16: Data samples from the datasets we use to evaluate UE methods: TriviaQA, NaturalQA, WebQA, and GSM8K.
@@ Appendix D Experimental Details
Datasets. To train the LARS model, for each TriviaQA and NaturalQA training split, we randomly select $\sim$13k samples resulting in $\sim$60k sampled unique QA pairs. We use all of the train split of GSM8K containing $\sim$8k samples. To evaluate the UE methods we use 4 datasets: $\sim$9k samples from the TriviaQA validation split, the validation set of NaturalQA consisting of $\sim$3500 samples, $\sim$6k samples coming from the train and validation sets of WebQA combined, and complete test split of GSM8K containing $\sim$1k samples.
@@
Computational Cost. We use 40 GB Nvidia A-100 GPUs for all the experiments. The total GPU-hours for training a LARS model with a calibration dataset generated from $\sim$13k questions is approximately 4. Labeling of the calibration data for one dataset and one model takes approximately 30 GPU-hours. Getting all the results in Tables 6 and 8 compromises $\sim$300 GPU-hours excluding LARS training. All presented results are obtained with a single run.
@@
User: You are a helpful, respectful
and honest assistant. Give precise,
short, one sentence answers to
given questions. Question:{question}
@@
The prompt for the teacher models explained in Section C.4 is as follows: